OpenStudy (anonymous):

–7 • (–11) A. 18 B. –77 C. 77 D. –18 i need help

OpenStudy (anonymous):

it's c

jhonyy9 (jhonyy9):

-7 *(-11) = 77

jhonyy9 (jhonyy9):

c.

OpenStudy (anonymous):

Learn those formulas and it will be easy for you, good luck :) (-) * (-)=+ (+) * (+)=+ (-) * (+)=- (+) * (-)=-

OpenStudy (anonymous):

thank you